# Compute-Constrained Data Selection

### v2

@@ Abstract
Data selection can reduce the amount of training data needed to finetune LLMs; however, the efficacy of data selection scales directly with its compute. Motivated by the practical challenge of compute-constrained finetuning, we consider the setting in which both the cost of selecting data and training are budgeted for. We first formalize the problem of data selection with a cost-aware utility function, and model the data selection problem as trading off initial-selection cost for training gain. We run a comprehensive sweep of experiments across multiple tasks, varying compute budget by scaling finetuning tokens, model sizes, and data selection compute. These experiments show the validity of this model in real-world experiments. Interestingly we find that many powerful data selection methods are almost never compute-optimal, and that cheaper data selection alternatives dominate both from a theoretical and empirical perspective.
@@ 1 Introduction
The growth of large language models (LLMs) has motivated research into their resource profiles. The compute cost of training LLMs is substantial, and, in practice, the total compute budget is predetermined: the number of accelerators and their usage hours are allocated in advanced. Thus, it is critical to determine the optimal allocation of resources under a budget constraint. Past work on compute-optimal LLMs (Hoffmann et al., 2022) studies if one could attain a better perplexity for a given pretraining compute budget by balancing architecture and training decisions.
Similar resource questions exist during post-training finetuning of LLMs. We assume a common setting where a single-base LLM needs to be finetuned for a downstream task. Numerous works have sought to induce certain abilities by training from large instruction tuning datasets (Sanh et al., 2021; Wei et al., 2021; Mishra et al., 2021). There are several different resource constraints which might make this challenging. For example work on parameter efficient fine-tuning, targets improving the memory-usage of this stage (Hu et al., 2021). In this work we focus instead on compute-constrained finetuning.
A promising approach to reducing compute requirements for finetuning is data selection. Data selection is a foundational approach in machine learning where the objective is to create an minimal dataset from a collection of data (John, 1975). Given the large computational cost of each gradient step, reducing dataset size is an appealing way to reduce this resource usage. Moreover, as larger and more diverse instruction tuning collections become available, likely only a subset of the data provides the most value for any given task. Recent work has shown that careful data selection can vastly increase the effectiveness of finetuning per step (Chen et al., 2023; Zhou et al., 2024).
@@
The results from this study argue that complex data selection methods are almost never Pareto-optimal in the compute-constrained setting and that simple statistical methods such as sparse retrieval should be preferred. In particular, powerful data selection methods that use model perplexity or gradient information tend to be FLOP inefficient both from a theoretical and empirical perspective. That is not to say these methods are not performance effective in improving the effectiveness of finetuning. For example they should be used in settings with repeated training with different tasks on the same underlying models. We additionally fit parameteric models to quantify the effectiveness of various approaches in a compute-aware manner. We hope that this framework and setting can motivate further research into cheaper data selection methods that can produce better models with less compute. Codebase and datasets to reproduce our results are available at https://github.com/oseyosey/CCDS.
@@ 2 Related Work
Data Selection for Language Models. Data selection takes the full training data as input and chooses a subset to the train (Albalak et al., 2024). It can be viewed as a coreset selection problem (Mirzasoleiman et al., 2020; Killamsetty et al., 2021b), which aims to select a subset from the given training dataset such that the model achieves performance similar to the full dataset.
@@
Task-Specific Finetuning from General-Purpose Instruction Datasets. While data selection can be applied in any finetuning setting, it is most impactful as a method to train a targeted model from a general-purpose dataset. In the case of LLMs, this setting is commonly training a task-specific model from an “instruction-tuning” dataset. Several general-purpose instruction tuning datasets exist for LLM post-training (Sanh et al., 2021; Wei et al., 2021; Mishra et al., 2021; Chung et al., 2024). Instruction-tuned models can handle a variety of possible inputs for downstream use cases as either classification or generative model (Wang et al., 2023b; a).
Instruction-tuned models are the most effective model for many downstream tasks; however they require training on a very large and expensive set of data. To get around this issue, models have demonstrated strong results using small subsets of instruction tuning data (Chen et al., 2023; Lu et al., 2023; Zhou et al., 2024). As the size of dataset increases, automated measures of quality selection has become a growing focus, particularly when many targeted models are needed. Therefore while instruction-tuning is not the direct focus of this work, it provide a real-world applications of compute-constrained data selection.
@@ 3 Background
The goal of data selection is to choose a subset of data points from a large dataset to optimize model performance on a target task. In a learning task, we are given a large training set, $\mathcal{D}$, a target test dataset, $\mathcal{T}$, and a validation set, $\mathcal{V}$. Our goal is to find the optimal subset $\mathcal{S}\subseteq\mathcal{X}$ such that the model $\theta=T(\mathcal{S})$ trained on $\mathcal{S}$ maximizes the performance on $\mathcal{T}$ under a given data constraint:
@@
This problem is challenging to solve in the general case, particularly without access to the test set $\mathcal{T}$. Approaches to the problem therefore commonly make two implicit assumptions: (1) the performance function, $P(\mathcal{T};T(\mathcal{S}))$, is monotonic and submodular, non-increasing marginal utility, in the dataset chosen (Kirchhoff & Bilmes, 2014), and (2) the validation set $\mathcal{V}$ correlates with the test set $\mathcal{T}$. Under these assumptions, we can argue for a greedy data selection approach (Kirchhoff & Bilmes, 2014). This allows decomposing the total objective, $P(\mathcal{T};T({\mathcal{S}}))$, by considering the contribution of individual training points to the performance on the validation set $P(\mathcal{V};T(\{x\}))$ for each $x\in\mathcal{S}$.
To estimate the marginal contribution of $x\in\mathcal{S}$, most data selection methods use a utility function $v(x;\mathcal{V})$—as a proxy to $P(\mathcal{V};T(\{x\}))$—to give the utility of each data point $x$ based on its relevance (Albalak et al., 2024). By ranking the data points $\mathcal{D}$ based on $v$ and selecting those that minimize the total utility within the data budget $K$, greedy data selection aims to obtain a high-performing subset $\mathcal{S}^{*}$. To summarize, we consider data selection methods that target Equation 1 with a two-step greedy algorithm: score all points and then select points up to the budget $K$.
@@ 4 Compute-Constrained Data Selection
Assuming we at minimal require the computation of a utility function over the dataset, we can define the compute-constrained data selection objective as,
@@ 4.1 Compute Cost of Data Selection Utilities
Method Utility Function Computational Cost $C_{\text{forward}}(x)$
@@
To make these costs more tangible, we consider four classes of data selection in this work, that represent three different levels of compute. This section summarizes their main properties, i.e. their core utility functions and computational costs. Table 1 describes the important computational costs from this section.
@@
Lexicon data selection methods utilize statistical properties of text to evaluate the relevance of data points without relying on deep learning models. One of the most effective lexicon-based methods is BM25 (Robertson et al., 2009; Silva & Barbosa, 2024), which scores data points based on the frequency of terms. The utility function $v_{\text{BM25}}(x;\mathcal{V})$ assigns a relevance score to each data point by averaging the BM25 scores with the validation, $v_{\text{BM25}}(x)=\frac{1}{|\mathcal{V}|}\sum_{x^{\prime}\in\mathcal{V}}\text{BM25}(x,x^{\prime})$. Since the algorithm can be run efficiently with a single-core cpu, the data selection FLOPs are almost 0.
@@
Different implementation of these general categories leads to different cost. A detailed cost analysis of our selected data selection methods can be found in the Appendix B.
@@ 5 Modeling the Compute-Performance Relationship
Figure 1: Simulation of Performance under Constraints. $P(k)=\bar{P}\times\left(1-\exp\left(-\lambda\frac{C(k)}{C(|\mathcal{D}|)}\right)\right)$ The behavior of different data selection methods using our performance model. (Left-Small Budget) The Lexicon method may consistently outperform more advanced data selection methods if their initial cost is too high. Under our assumptions gradient can never be optimal as its cost exceeds 1 epoch of training. (Middle-Medium Budget) The perplexity method can become optimal once the total cost exceeds a given amount. (Right-Large Budget) The gradient methods can be optimal if training is more expensive than the fixed-cost, for example if using a much larger base model then data selection model. The simulation shows that the compute-optimal data selection method changes as a function of the compute budget and the performance rate associated with each method.
@@
Using methodology similar to Hoffmann et al. (2022), the parameter $\lambda$ can be fit on the empirical measurements, which yields data-driven estimates. See Appendix C for more details on derivations and the fitting procedure.
@@ 6 Experimental Setup
We follow Wang et al. (2023a) and curate a representative sample of instruction-tuned datasets as listed in Table 8. This includes: (1) datasets generated by researchers from existing NLP datasets, such as COT (Wei et al., 2022) and Flan V2 (Longpre et al., 2023) ; (2) datasets written by humans from scratch specifically for instruction tuning, including Dolly (Conover et al., 2023) and Open Assistant 1 (Köpf et al., 2024).
@@
For all experiments, we train transformer language models with Llama architecture and tokenizer (Touvron et al., 2023; Dubey et al., 2024). Models range from 7B parameters to 70B parameters, and are trained for close to 100 million total fine-tuning tokens. Our experiments closely follow prior work on training and evaluating instruction-tuned models (Wang et al., 2023a; Ivison et al., 2023).
@@
Figure 2: Performance for Different Data Selection Methods. We show all of our different runs for a given model size, where each scatter point is the final target task performance of a single run. (A, B, C) show MMLU results across three model sizes, while (D, E, F) present BBH results across three model sizes. For each run, we determine the optimal finetuning strategy—a combination of data selection method and number of finetuning tokens—that achieves the highest performance under a particular FLOPs budget. We fit a pareto front in dashed line based on these optimal strategies. At small and medium compute budgets (A, B, D, E), cheaper data selection methods like BM25 and EMBED outperform PPL and LESS, which rely on model information. At larger compute budgets (C, F), however, PPL and LESS become compute-optimal after using 5% of the fine-tuning tokens.
@@ 7 Results
Note that for PPL and Gradient, a 7B model of the same class is always used for data selection; whereas for Embed a small encoder model is used. For MMLU, PPL is implemented as Mid-PPL, where points in the middle of perplexity score distribution is being ranked first. For BBH, PPL is implemented as Top-PPL, where points in the top of perplexity score distribution is ranked first.
@@
Figure 3: Parametric Fit of Performance with Compute-Constrained Data Selection. We fit a parametric model of the performance in Equation 3 and display that as curves to pair with the empirical results as scatter points. (A, B, C) show MMLU results and their parametric fit across three model sizes, while (D, E, F) present BBH results and their parametric fit across three model sizes. For each fixed model sizes, we show the finetuned pareto frontier in dashed gray line, which is a line in the log-linear space. Note that the x-axis is log-scale.
Fit of Compute-Performace Relationship. In Section 5 we propose a parameteric model for the relationship of data selection compute to model performance, Equation 3. This formula assumed that data selection diminish in utility and that performance is a direct function of compute. To verify these assumption, we fit the parametic form to the the empirical curves by fitting the $\lambda$ parameter and allowing tolerance in $P_{0},\bar{P}$, per method and dataset. We model all final performances from our experiments as a parametric function of the compute budget $P(k)$.
Figure 3 shows fittings for each curve. Since the shape of these curves is dataset and model dependent, we are unable to forward predict specific performances from additional data selection compute. However the close fit of the curves indicates that the underlying model — significant performance gain at the cost of exponentially more compute investment — holds up in practical settings.
@@ 8 Analysis
Another interesting setting is when the goal is to train multiple models from the same large training set, where each targets a different task. In this setting, the gradient information calculation can be amortized between target tasks. However, even in this setting the cost of gradient calculation is still quite high. As shown in Figure 4, with Llama2 7B as the data selection model, LESS needs 20 tasks to surpass the finetuned pareto frontier. We perform similar analysis for other model sizes in Appendix G.
@@
Figure 6 presents the Jaccard similarity of data selected by different data selection methods. We observe that BM25 and Embed tend to select highly similar data, which explains their comparable performance. In contrast, the data selected by LESS bears the least resemblance to the other methods. We perform similar analysis for different target tasks in Appendix H.
@@ B.4 LESS
Appendix C Fitting Parametric Function

### v4

@@ Abstract
Data selection can reduce the amount of training data needed to finetune LLMs; however, the efficacy of data selection scales directly with its compute. Motivated by the practical challenge of compute-constrained finetuning, we consider the setting in which both the cost of selecting data and training are budgeted for. We first formalize the problem of data selection with a cost-aware utility function, and model the data selection problem as trading off initial-selection cost for training gain. We run a comprehensive sweep of experiments across multiple tasks, varying compute budget by scaling finetuning tokens, model sizes, and data selection compute. Interestingly we find that many powerful data selection methods are almost never compute-optimal, and that cheaper data selection alternatives dominate both from a theoretical and empirical perspective. For compute-optimal training, we find that perplexity and gradient data selection require training-to-selection model size ratios of 5x and 10x, respectively.
@@ 1 Introduction
The growth of large language models (LLMs) has motivated research into their resource profiles. The compute cost of training LLMs is substantial, and, in many cases, the total compute budget is predetermined: the number of accelerators and their usage hours are allocated in advanced. Thus, it is critical to determine the optimal allocation of resources under a budget constraint. Past work on compute-optimal LLMs (Hoffmann et al., 2022) studies if one could attain a better perplexity for a given pretraining compute budget by balancing architecture and training decisions.
Similar resource questions exist during post-training finetuning of LLMs. We assume a common setting where a single-base LLM needs to be finetuned for a downstream task. Numerous works have sought to induce certain abilities by training from large instruction tuning datasets (Sanh et al., 2021; Wei et al., 2021; Mishra et al., 2021). There are several different resource constraints which might make this challenging. For example, parameter-efficient fine-tuning methods like LoRA (Hu et al., 2021) aim to reduce memory usage during finetuning by updating only a small subset of the model’s parameters. In this work we focus instead on compute-constrained finetuning.
A promising approach to reducing compute requirements for finetuning is data selection. Data selection is a foundational approach in machine learning where the objective is to create a minimal dataset from a collection of data (Hart, 1968; John, 1975). Given the large computational cost of each gradient step, reducing dataset size is an appealing way to reduce this resource usage. Moreover, as larger and more diverse instruction tuning collections become available, likely only a subset of the data provides the most value for any given task. Recent work has shown that careful data selection can vastly increase the effectiveness of finetuning per step (Chen et al., 2023; Zhou et al., 2024).
@@
The results from this study argue that complex data selection methods are almost never Pareto-optimal in the compute-constrained setting and that simple statistical methods such as sparse retrieval should be preferred. In particular, powerful data selection methods that use model perplexity or gradient information tend to be FLOP inefficient both from a theoretical and empirical perspective. That is not to say these methods are ineffective; for example, they should be used in settings with repeated training with different tasks on the same underlying models. We fit parameteric models to quantify the effectiveness of various approaches in a compute-aware manner. Using these fits, we extrapolate and empirically validate that, for comptue-optimal finetuning, perplexity and gradient data selection require training-to-selection model size ratios of 5x and 10x, respectively.
We hope that this framework and setting can motivate further research into cheaper data selection methods that can produce better models with less compute. Codebase and datasets to reproduce our results are available at https://github.com/oseyosey/CCDS.
@@ 2 Related Work
Data Selection for Language Models. Data selection takes the full training data as input and chooses a subset to train (Albalak et al., 2024). It can be viewed as a coreset selection problem (Mirzasoleiman et al., 2020; Killamsetty et al., 2021b), which aims to select a subset from the given training dataset such that the model achieves performance similar to the full dataset.
@@
Task-Specific Finetuning from General-Purpose Instruction Datasets. While data selection can be applied in any finetuning setting, it is most impactful as a method to train a targeted model from a general-purpose dataset. In the case of LLMs, this setting is commonly training a task-specific model from an “instruction-tuning” dataset (Sanh et al., 2021; Wei et al., 2021; Mishra et al., 2021).
Instruction-tuned models are effective for many downstream tasks; however they require training on a very large and expensive set of data. To get around this issue, models have demonstrated strong results using small subsets of instruction tuning data (Chen et al., 2023; Lu et al., 2023; Zhou et al., 2024). As datasets grow, automated measures of quality selection has become a growing focus, particularly when many targeted models are needed. Therefore while instruction-tuning is not the direct focus of this work, it provide a real-world applications of compute-constrained data selection.
@@ 3 Background
The goal of data selection is to choose a subset of data points from a large dataset to optimize model performance on a target task. In a learning task, we are given a large training set, $\mathcal{D}$, a target test dataset, $\mathcal{T}$, and a validation set, $\mathcal{V}$. Our goal is to find the optimal subset $\mathcal{S}\subseteq\mathcal{D}$ such that the model $\theta=T(\mathcal{S})$ trained on $\mathcal{S}$ maximizes the performance on $\mathcal{T}$ under a given data constraint:
@@
This problem is challenging to solve in the general case, particularly without access to the test set $\mathcal{T}$. Approaches to the problem therefore commonly make two implicit assumptions: (1) the performance function, $P(\mathcal{T};T(\mathcal{S}))$, is monotonic and submodular, non-increasing marginal utility, in the dataset chosen (Kirchhoff & Bilmes, 2014), and (2) the validation set $\mathcal{V}$ is IID with the test set $\mathcal{T}$. Under these assumptions, we can argue for a greedy data selection approach (Kirchhoff & Bilmes, 2014). This allows decomposing the total objective, $P(\mathcal{T};T({\mathcal{S}}))$, by considering the contribution of individual training points to the performance on the validation set $P(\mathcal{V};T(\{x\}))$ for each $x\in\mathcal{S}$.
To estimate the marginal contribution of $x\in\mathcal{S}$, most data selection methods use a utility function $v(x;\mathcal{V})$—as a proxy to $P(\mathcal{V};T(\{x\}))$—to give the utility of each data point $x$ based on its relevance (Albalak et al., 2024). By ranking the data points $\mathcal{D}$ based on $v$ and selecting those that maximize the total utility within the data budget $K$, greedy data selection aims to obtain a high-performing subset $\mathcal{S}^{*}$. To summarize, we consider data selection methods that target Equation 1 with a two-step greedy algorithm: score all points and then select points up to the budget $K$.
@@ 4 Compute-Constrained Data Selection
Assuming we at minimum require the computation of a utility function over the dataset, we can define the compute-constrained data selection objective as
$\mathcal{S}^{*}=\underset{\mathcal{S}\subseteq\mathcal{D}}{\arg\max}\quad P(\mathcal{V};T(\mathcal{S}))$ (2)
@@ 4.1 Compute Cost of Data Selection Utilities
Method Utility Function Computational Cost $\mathbf{C_{\text{{forward}}}(x)}$
@@
To make these costs more tangible, we consider four classes of data selection in this work, that represent three different levels of compute. This section summarizes their main properties, i.e. their core utility functions and computational costs.
@@
Lexicon data selection methods utilize statistical properties of text to evaluate the relevance of data points without relying on deep learning models. One of the most effective lexicon-based methods is BM25 (Robertson et al., 2009; Silva & Barbosa, 2024), which scores data points based on the frequency of terms. The utility function $v_{\text{BM25}}(x;\mathcal{V})$ assigns a relevance score to each data point by averaging the BM25 scores with the validation, $v_{\text{BM25}}(x)=\frac{1}{|\mathcal{V}|}\sum_{x^{\prime}\in\mathcal{V}}\text{BM25}(x,x^{\prime})$. Since the algorithm can be run with a single-core cpu, the data selection FLOPs are almost 0.
@@
While lexicon and embedding-based methods aim to select training samples similar to validation samples, perplexity and gradient-based methods focus on optimizing their effect on model loss. Different implementation of these general categories leads to different cost. A detailed cost analysis of our selected data selection methods can be found in the Appendix B.
@@ 5 Modeling the Compute-Performance Relationship
Figure 1: Simulation of Performance under Constraints. $P(k)=\bar{P}\times\left(1-\exp\left(-\lambda\frac{C(k)}{C(|\mathcal{D}|)}\right)\right)$ The behavior of different data selection methods using our performance model. (Left-Small Budget) The Lexicon method may consistently outperform more advanced data selection methods if their initial cost is too high. Under our assumptions gradient can never be optimal as its cost exceeds 1 epoch of training. (Middle-Medium Budget) The perplexity method can become optimal once the total cost exceeds a given amount. (Right-Large Budget) The gradient methods can be optimal if training is more expensive than the fixed-cost, for example if using a much larger base model than data selection model. The simulation shows that the compute-optimal data selection method changes as a function of the compute budget and the performance rate associated with each method.
@@
Using methodology similar to Hoffmann et al. (2022), the parameter $\lambda$ can be fit on the empirical measurements, which yields data-driven estimates. See Section C.1 for more details on derivations and the fitting procedure.
@@ 6 Experimental Setup
We follow Wang et al. (2023) and curate a representative sample of instruction-tuned datasets as listed in Table 8. This includes: (1) datasets generated by researchers from existing NLP datasets, such as COT (Wei et al., 2022) and Flan V2 (Longpre et al., 2023) ; (2) datasets written by humans from scratch specifically for instruction tuning, including Dolly (Conover et al., 2023) and Open Assistant 1 (Köpf et al., 2024).
@@
For all experiments, we train transformer language models with Llama architecture and tokenizer (Touvron et al., 2023; Dubey et al., 2024). Models range from 7B parameters to 70B parameters, and are trained for close to 100 million total fine-tuning tokens. Our experiments closely follow prior work on training and evaluating instruction-tuned models (Wang et al., 2023; Ivison et al., 2023).
@@
Figure 2: Performance for Different Data Selection Methods. We show all of our different runs for a given model size, where each scatter point is the final target task performance of a single run. (A, B, C) show MMLU results across three model sizes, while (D, E, F) present BBH results across three model sizes. For each run, we determine the optimal finetuning strategy—a combination of data selection method and number of finetuning tokens—that achieves the highest performance under a particular FLOPs budget. We fit a pareto front in dashed line based on these optimal strategies, which is a line in the linear-log space. At small and medium compute budgets (A, B, D, E), cheaper data selection methods like BM25 and EMBED outperform PPL and LESS, which rely on model information. At larger compute budgets (C, F), however, PPL and LESS become compute-optimal after using 5% of the fine-tuning tokens.
@@ 7 Results
Note that for PPL and Gradient, a 7B model of the same model family is always used for data selection; whereas for Embed a small encoder model is used. For MMLU, PPL is implemented as Mid-PPL. For BBH, PPL is implemented as Top-PPL.
The fine-tuned efficient Pareto frontier comprises all runs that are Pareto-optimal with respect to compute (x-axis) and performance (y-axis). These runs represent the most efficient choices, providing the best possible performance for a given compute budget under specific data selection methods and training token lengths. Furthermore, we model the efficient computational frontier using a power law, specifically of the form $P(C)=a\log(C)+b,$ where $a$ and $b$ are parameters fitted to the data. The fitted function is depicted as a dashed gray line in the figures.
@@
Figure 3: Parametric Fit of Performance with Compute-Constrained Data Selection. We fit a parametric model of the performance in Equation 3 and display that as curves to pair with the empirical results as scatter points. (A, B, C) show MMLU results and their parametric fit across three model sizes, while (D, E, F) present BBH results and their parametric fit across three model sizes.
Fit of Compute-Performance Relationship. In Section 5 we propose a parameteric model for the relationship of data selection compute to model performance, Equation 3. This formula assumed that data selection diminish in utility and that performance is a direct function of compute. To verify these assumption, we fit the parametic form to the the empirical curves by fitting the $\lambda$ parameter and allowing tolerance in $P_{0},\bar{P}$, per method and dataset. We model all final performances from our experiments as a parametric function of the compute budget $P(k)$.
Extrapolation from Parametric Fits. Figure 3 shows the fitted curves for each method. Since the shape of these curves is dataset and model dependent, we cannot predict exact performance for different levels of data selection compute. However, the close fit of the parametric models obtained from smaller models allows us to estimate the compute-optimal ratio between the training model size and the selection model size. For perplexity-based data selection, our extrapolation suggests that the method becomes compute-optimal when the training model is 5x larger than the data selection model—around 35B parameters. For gradient-based data selection, our extrapolation indicates that the training model needs to be approximately 10x larger than the data selection model to be compute-optimal—around 70B parameters. Appendix G presents our results and details the extrapolation.
@@ 8 Analysis
Another interesting setting is when the goal is to train multiple models from the same large training set, where each targets a different task. In this setting, the gradient information calculation can be amortized between target tasks. However, even in this setting the cost of gradient calculation is still quite high. As shown in Figure 4, with Llama2 7B as the data selection model, LESS needs 20 tasks to surpass the finetuned pareto frontier. We perform similar analysis for other model sizes in Appendix H.
@@
Figure 6 presents the Jaccard similarity of data selected by different data selection methods. We observe that BM25 and Embed tend to select highly similar data, which explains their comparable performance. In contrast, the data selected by LESS bears the least resemblance to the other methods. We perform similar analysis for different target tasks in Appendix I.
@@ B.4 LESS
Appendix C Parametric Function
C.1 Fitting Parametric Function